# Benchmark of segmentation techniques for pelvic fracture in CT and X-ray: Summary of the PENGWIN 2024 challenge

### v1

@@ 2.2.2 Fracture segmentation in X-ray
Early applications of deep learning for fracture analysis in X-ray imaging primarily focused on fracture detection and classification, with some approaches incorporating bone segmentation as a pre-processing step (Joshi and Singh, 2020).
@@
Fracture detection can be performed at the image level, where a model determines the presence or absence of fractures. For instance, Cheng et al. (2019) developed a convolutional network to detect hip fractures from plain frontal pelvic radiographs, achieving an accuracy of 91%. To enhance interpretability, they further applied gradient-weighted class activation mapping (Grad-CAM) to visualize fracture regions, yielding an accuracy of 95.9% for lesion identification. Similar binary fracture classification approaches have been studied on other anatomies like long bones and spine (Wang et al., 2022b).
@@
Beyond image-level classification, fractures can be localized more precisely using bounding boxes or segmentation contours. This is particularly valuable for hairline or incomplete fractures, where the fracture line is clearly visible in the radiograph, though the entire fragment cannot be isolated. Traditional image processing and machine learning-based methods have been explored for fracture detection using bounding box representations (Myint et al., 2018; Hržić et al., 2019; Hardalaç et al., 2022). For example, Joshi et al. (2022) employed a feature pyramid network with transfer learning to perform both fracture detection (bounding boxes) and segmentation (contours) in wrist bone fractures. Turk et al. (2024) investigated the effect of different backbone architectures on fracture segmentation in X-ray images. Their study utilized a dataset of 4,083 scans covering hand, leg, hip, and shoulder fractures, with 4,497 distinct fracture sites. The highest-performing method achieved an Intersection over Union (IoU) exceeding 0.93.
@@
Additionally, in the pelvic region, a recent study proposed the use of deep networks to segment the entire pelvis in X-ray images in the presence of fractures, demonstrating the potential of deep learning methods in handling fracture cases (Lee et al., 2024). In spinal X-ray imaging, efforts have also been made to detect, segment, and assess vertebral compression fractures. Since spinal fractures are often non-displaced and lack clearly visible fracture lines, existing segmentation methods have primarily focused on extracting the full outline of the affected vertebrae (Kim et al., 2021).
@@
In short, similar to the related challenges discussed in Sec. 2.1, existing studies on X-ray fracture analysis have prioritized fracture detection, with some leveraging deep learning-based segmentation as a means to refine localization. While segmentation provides a more precise localization than bounding boxes to support trauma diagnosis, current methods lack the ability to delineate fracture lines accurately or extract individual fractured fragments. Bone fragment segmentation in X-ray images remains particularly challenging due to fragment overlap—especially in the pelvis, where the ring structure complicates separation—and the presence of soft tissues in the projection image, which obscures boundaries and limits the effectiveness of standard segmentation approaches.
3 Challenge description
3.1 Challenge setup
@@
The PENGWIN Challenge 2024 is a one-time event with fixed submission deadline comprising two independent tasks, with participants free to compete in either or both. Each task included a development phase and a final test phase, following a structured timeline. The training data was released on April 5, 2024, enabling participants to develop and refine their models. The validation phase, beginning on May 15, 2024, allowed teams to test their models on a separate validation set multiple times, with the results displayed on a public leaderboard. In the final test phase, participants submitted their models as Docker containers on the Grand Challenge platform [^1] Challenge website: https://pengwin.grand-challenge.org/ by August 19, 2024. The final results were announced on August 31, 2024, and the top three teams in each task were awarded cash prizes and certificates, conditional upon the open-sourcing of their code in a public repository. The top three teams were invited to present their methods at MICCAI 2024 in October. Only the members of the organizing team have access to the test cases. Members of the organizers’ institutes could participate but were not eligible for awards. Participating teams could publish their own results separately, with a one-year embargo period.
@@
3.2 Data
3.2.1 Curation of the CT dataset
The CT data were collected from 150 patients who were planning to undergo pelvic reduction surgery between 2017 and 2023 from six medical centers in China, including Beijing Jishuitan Hospital (JST), Foshan Hospital of TCM (FSHTCM), the First Bethune Hospital of Jilin University (JLUFH), the Third Bethune Hospital of Jilin University (JLUTH), Nanfang Hospital (NFH), and Tianjin Hospital (TJH). The CT scans were acquired from a diverse array of machines symmarized in Table 2. The retrospective data usage has been approved by the Ethics Committee of the institutes.
Table 2: List of CT scanners and number of cases in different centers.
@@ 3.2.1 Curation of the CT dataset
The dataset includes patients aged 16 to 94 (44.1 ± 16.8) years, with a gender distribution of 63 females and 87 males. The dataset encompasses a variety of fracture types, including pelvic ring dislocation (PRD, 5 cases), unilateral hip fracture (UHF, 54 cases), bilateral hip fractures (BHF, 31 cases), sacral fracture (SF, 5 cases), and sacral and hip fractures (SHF, 55 cases). The number of pelvic bone fragments per case ranged from 3 to 12 (5.7 ± 1.4).
The average voxel spacing is (0.83, 0.83, 0.89) mm. In cases where extra anatomies like abdomen or femoral shaft were included, the volumes were cropped to contain the pelvic region. The image dimensions average at (488, 426, 323). All the scans were de-indentified and converted to MetaImage Medical (MHA) format.
The 150 cases are split into 100 for training, 20 for validation, and 30 for testing. To ensure a balanced representation in the test set, we employed stratified random sampling based on the five primary fracture types, so that the training and test sets have the same fracture type distribution.
3.2.2 Annotation
The dataset is processed by two experienced annotators each with more than five years of medical imaging experience and a senior expert with more than 15 years of orthopedic surgical planning experience. The data annotation was structured into a four-step workflow:
@@ 3.2.2 Annotation
    Initial automatic segmentation: We employ a pre-trained segmentation network based on the nnUNet framework to produce preliminary segmentations for the hipbones and sacrum. This network was trained on roughly 1,000 pelvic CT images in the CTPelvic1K dataset, with the majority of them not presenting any fractures (Liu et al., 2021).
@@
    Manual refinement of anatomical labels: The initial anatomical labels undergo a meticulous refinement process by annotators.
@@
    Identification of fractured fragments: Leveraging the refined anatomical labels, the annotators identify and label fractured bone fragments.
@@
    Expert validation: As a final checkpoint, a senior expert rigorously reviews, modifies, and authenticates the annotated labels, ensuring their precision and consistency.
In steps 2 and 3, the annotators were instructed to annotate three anatomical labels—left hipbone, right hipbone, and sacrum—as well as the bone fragments within each bone using the 3D Slicer software. If a bone is intact and does not have any fractures, the only fragment corresponds to its anatomical mask. Fragments with a volume smaller than 500 mm3 were omitted. The ground-truth labels were also saved in MHA format.
@@
Fig. 1 illustrates an example CT scan along with its annotated segmentation mask. In this instance segmentation task, the number of fragment instances varies across cases. The label indexing follows a two-level hierarchical structure: first, fragments are classified based on their anatomical origin, and second, they are ordered by fragment size. The CT training set is publicly available in an open repository under a CC BY-NC-SA license [^2] PENGWIN CT dataset: https://doi.org/10.5281/zenodo.10927452.
3.2.3 Generation of the X-ray dataset
Figure 2: Example simulated X-ray image and the annotated segmentation.
The PENGWIN X-ray dataset comprises digitally reconstructed radiographs (DRRs) that simulate intraoperative C-arm X-ray images, including pelvic anatomy, fracture fragments, and surgical tools. To ensure robustness against surgical hardware interference, instrument poses were randomly varied across images. The X-ray images were synthesized from CT data using the DeepDRR framework, which simulates realistic X-ray projections through a combination of analytic ray-tracing and learning-based scatter estimation and noise injection techniques (Unberath et al., 2018). The generation pipeline involves the following steps:
  1. 1.
    Material decomposition is applied to the CT volume to distinguish air, soft tissue, and bone.
  2. 2.
    Forward projection is conducted using material- and spectrum-aware ray-tracing, simulating X-ray attenuation along each ray path.
  3. 3.
    Scatter estimation is performed using a deep learning-based convolutional network trained on Monte Carlo simulation data to predict scatter contributions across the projection image.
  4. 4.
    Noise injection is applied, incorporating quantum noise and electronic readout noise to approximate realistic X-ray image formation.
@@ 3.2.3 Generation of the X-ray dataset
The simulated X-ray dataset was generated from the 150 CT scans described above, following the same data partitioning scheme. For the training and validation sets, 400 X-ray images were simulated per CT scan, while for the test set, 20 X-ray images were generated per CT scan, resulting in a total of 600 test cases. The images were provided in binary Tag Image File Format (TIFF), with multi-label segmentation masks encoded as single-channel binary TIFF images to comply with the challenge platform requirements. Additionally, the encoding and decoding scripts for handling segmentation files were made available to participants to ensure consistency in data processing. The X-ray training set is publicly available in an open repository under a CC BY-NC-SA license [^3] PENGWIN X-ray dataset: https://doi.org/10.5281/zenodo.10913195.
3.3 Assessment method
3.3.1 Evaluation metrics
@@
For both the CT task and the X-ray task, the primary evaluation metrics included intersection over union (IoU), the 95th percentile Hausdorff distance (HD95), and average symmetric surface distance (ASSD). IoU was computed on solid masks, while HD95 and ASSD were computed on surface contours. The adoption of these metrics aligns with the recommendations from the metrics reloader website and has been validated in the study by Faghani et al. (2022).
IoU measures the overlap between the predicted segmentation mask $S_{p}$ and the ground truth mask $S_{g}$, and is defined as:
$\text{IoU}=\frac{|S_{p}\cap S_{g}|}{|S_{p}\cup S_{g}|},$ (1)
where $\cap$ denotes the intersection and $\cup$ denotes the union of the two masks.
HD95 quantifies the worst-case segmentation error by measuring the 95th percentile of the distances between the predicted and ground-truth segmentation contours. It is defined as:
@@
$\text{ASSD}=\frac{1}{|C_{p}|+|C_{g}|}\left(\sum_{x\in C_{p}}\inf_{y\in C_{g}}d(x,y)+\sum_{y\in C_{g}}\inf_{x\in C_{p}}d(x,y)\right),$ (3)
where distances are computed symmetrically in both directions.
@@
3.3.2 Ranking scheme
@@
    Metric calculation: For each individual test case, evaluation metrics were computed for both fracture segmentation and anatomical segmentation.
@@
    Tie-breaking: In the event of a tie in the final ranking, the container execution time of the submission was used as a tie-breaker, with faster models receiving a higher rank.
4 Results
4.1 Participation
@@ 4.2 Algorithm summary
4.2.1 CT segmentation
Table 3: Summary of the top-performing participating algorithms for CT segmentation. Spatial augmentation includes random scaling, mirroring, and flipping. Intensity augmentation includes noising, blurring, and histogram adjustments.

### v2

@@ II-B2 Fracture Segmentation in X-Ray
Early applications of deep learning for fracture analysis in X-ray primarily focused on fracture detection and classification, with some approaches incorporating bone segmentation as a pre-processing step [joshi2020survey].
@@
Fracture detection can be performed at the image level, where a model determines the presence or absence of fractures. For instance, Cheng et al. [cheng2019application] developed a convolutional network to detect hip fractures from plain frontal pelvic radiographs. To enhance interpretability, they further applied gradient-weighted class activation mapping (Grad-CAM) to visualize fracture regions. Similar binary fracture classification approaches have been studied on other anatomies like long bones and spine [wang2022osteoporotic].
@@
Beyond image-level classification, fractures can be localized more precisely using bounding boxes or segmentation contours. This is particularly valuable for hairline or incomplete fractures, where the fracture line is clearly visible in the radiograph, though the entire fragment cannot be isolated. Traditional image processing and machine learning-based methods have been explored for fracture detection using bounding box representations [myint2018analysis, hrvzic2019local, hardalacc2022fracture]. For example, Joshi et al. [joshi2022deep] employed a feature pyramid network to perform both fracture detection (bounding boxes) and segmentation (contours) in wrist bone fractures. Turk et al. [turk2024impact] investigated the effect of different backbone architectures on fracture segmentation in X-rays. Their study utilized 4,083 images covering hand, leg, hip, and shoulder fractures.
@@
Additionally, in the pelvic region, a recent study proposed the use of deep networks to segment the entire pelvis in X-ray in the presence of fractures, demonstrating the potential of deep learning methods in handling fracture cases [lee2024deep]. In spinal X-ray imaging, efforts have also been made to detect, segment, and assess vertebral compression fractures. Since spinal fractures are often non-displaced and lack clearly visible fracture lines, existing methods have primarily focused on extracting the full outline of the affected vertebrae [kim2021automatic].
@@
In short, similar to the related challenges discussed in Sec. II-A, existing studies on X-ray fracture analysis have prioritized fracture detection, with some leveraging deep learning-based segmentation as a means to refine localization. While segmentation provides a more precise localization than bounding boxes to support trauma diagnosis, current methods lack the ability to delineate fracture lines accurately or extract individual fractured fragments.
III Challenge Description
III-A Challenge Setup
@@
The PENGWIN Challenge 2024 is a one-time event with fixed submission deadline comprising two tasks, with participants free to compete in either or both. Each task included a development phase and a final test phase. The training data was released on April 5, 2024, enabling participants to develop and refine their models. The validation phase, beginning on May 15, 2024, allowed teams to test their models on a separate validation set multiple times, with the results displayed on a public leaderboard. In the final test phase, participants submitted their models as Docker containers on the Grand Challenge platform [^1] Challenge website: https://pengwin.grand-challenge.org/. by August 19, 2024. The final results were announced on August 31, 2024, and the top three teams in each task were awarded cash prizes and certificates, conditional upon the open-sourcing of their code in a public repository. The top three teams were invited to present their methods at MICCAI 2024 in October. Only the organizing team have access to the test cases. Members of the organizers’ institutes could participate but were not eligible for awards.
@@
III-B Data
III-B1 Curation of the CT Dataset
The CT data were collected from 150 patients who were planning to undergo pelvic reduction surgery between 2017 and 2023 from six medical centers in China, with varying distributions (ranging from 6 to 58 scans per site) due to differences in data availability and institutional capacity. The CT scans were acquired from a diverse array of machines symmarized in Table I. The retrospective data usage has been approved by the Ethics Committee of the institutes.
TABLE I: List of CT scanners and number of cases in different centers.
@@ III-B1 Curation of the CT Dataset
The dataset includes patients aged 16 to 94 (44.1 ± 16.8) years, with a gender distribution of 63 females and 87 males. The dataset encompasses a variety of fracture types, including pelvic ring dislocation (PRD, i.e., dislocation of pubic symphysis and/or sacroiliac joint without fracture within single bone, 5 cases), unilateral hip fracture (UHF, 54 cases), bilateral hip fractures (BHF, 31 cases), sacral fracture (SF, 5 cases), and sacral and hip fractures (SHF, 55 cases). The number of pelvic bone fragments per case ranged from 3 to 12 (5.7 ± 1.4).
The average voxel spacing is (0.83, 0.83, 0.89) mm. In cases where extra anatomies like abdomen or femoral shaft were included, the volumes were cropped to contain the pelvic region. The image dimensions average at (488, 426, 323). All the scans were de-indentified and converted to MHA format. The 150 cases are split into 100 for training, 20 for validation, and 30 for testing. To ensure a balanced representation in the test set, we employed stratified random sampling based on the five primary fracture types, so that the training and test sets have the same fracture type distribution.
III-B2 Annotation
The dataset is processed independently by two experienced annotators each with more than five years of medical imaging experience and a senior expert with more than 15 years of orthopedic surgical planning experience. The annotators were instructed to annotate three anatomical labels—left hipbone, right hipbone, and sacrum—as well as the bone fragments within each bone using the 3D Slicer software. The data annotation was structured into a four-step workflow:
@@ III-B2 Annotation
    Initial automatic segmentation: We employ a pre-trained segmentation network based on the nnUNet framework to produce preliminary segmentations for the hipbones and sacrum. This network was trained on the CTPelvic1K dataset, with the majority of them not presenting any fractures [liu2021deep].
@@
    Manual refinement of anatomical labels: The initial anatomical segmentations were manually refined by trained annotators. This step involved removing any non-pelvic structures and restoring under-segmented regions, with particular attention paid to boundary regions near fracture surfaces where automatic segmentation is most prone to error. Refinement was performed using localized intensity thresholding and scissor-based region editing tools to ensure anatomically correct and topologically consistent bone surfaces.
@@
    Identification of fractured fragments: Leveraging the refined anatomical labels, the annotators delineate individual fracture fragments using scissor-based region separation and slice-by-slice manual painting tools to ensure accurate fragment boundaries.
@@
    Expert adjudication: The two independently produced annotations are then reviewed side-by-side by a senior expert, who selects the annotation that better reflects clinically realistic fracture separability. In cases where the two annotations are visually indistinguishable, selections are alternated across cases to avoid systematic bias. This adjudication step ensures label consistency while preserving clinical relevance.
Before annotation began, the annotators underwent a training process, which included standardized written instructions and reference examples covering common fracture patterns. A key aspect of the refinement procedure involves determining whether an incomplete fracture (a partially fractured but non-displaced cortical discontinuity) should be labeled as one fragment or two fragments. We provide the following rule-of-thumb to guide this decision: If the cortical separation is expected to cause independent movement of the two regions during surgical reduction, the bone is labeled as separate fragments; If the two regions are expected to move together as a single rigid unit, the region is labeled as one fragment. In addition, if a bone does not have any fracture, the only fragment corresponds to its anatomical mask. Fragments with a volume smaller than 500 mm3 were omitted. The ground-truth labels were also saved in MHA format.
@@
Fig. 1 illustrates an example CT scan along with its annotated segmentation mask. In this instance segmentation task, the number of fragment instances varies across cases. The label indexing follows a two-level hierarchical structure: first, fragments are classified based on their anatomical origin, and second, they are ordered by fragment size. The CT training set is publicly available under a CC BY-NC-SA license [^2] PENGWIN CT dataset: https://doi.org/10.5281/zenodo.10927452..
III-B3 Generation of the X-Ray Dataset
Figure 2: The X-ray simulation process of DeepDRR.
Figure 3: Example simulated X-ray image and the annotation.
The PENGWIN X-ray dataset comprises digitally reconstructed radiographs (DRRs) that simulate intraoperative C-arm X-ray images, including pelvic anatomy, fracture fragments, and surgical tools. To ensure robustness against surgical hardware interference, instrument poses were randomly varied across images. The X-ray images were synthesized from CT data using the DeepDRR framework, which simulates realistic projections through a combination of analytic ray-tracing and learning-based scatter estimation and noise injection techniques [unberath2018deepdrr]. As shown in Fig. 2, the generation pipeline involves the following steps:
First, material decomposition is applied. A 3D convolutional network classifies every voxel in the CT volume as air, soft tissue or bone. The resulting material map preserves each tissue’s true spatial distribution and provides the physical input needed for X-ray simulation.
Second, forward projection is conducted using material- and spectrum-aware ray-tracing. Using this map, DeepDRR traces virtual X-ray beams from the source to the detector. For each detector pixel $u$, the transmitted intensity is
$p(u)=\int_{E}p_{0}(E)\,\exp\!\Bigl(-\sum_{m\in M}(\mu/\rho)_{m}(E)\,\int_{\ell_{u}}\rho_{m}(x)\,d\ell\Bigr)\,dE,$ (1)
where $p_{0}(E)$ is the source spectrum, $(\mu/\rho)_{m}(E)$ the energy-dependent attenuation of material $m$, $\rho_{m}(x)$ its density along the ray $\ell_{u}$, and $M=\{\text{air},\ \text{soft tissue},\ \text{bone}\}$ is simulated materials. This step yields the primary, scatter-free projection.
Third, scatter estimation is performed. Real fluoroscopy contains low-frequency scatter that softens contrast. DeepDRR estimates this component with a lightweight ten-layer convolutional network trained on Monte-Carlo data. The predicted scatter image is added pixel-wise to the primary projection.
Finally, detector realism is achieved by adding two noise sources. Photon (quantum) noise is drawn from a Poisson distribution whose mean photon count,
$N(u)=\sum_{E}\frac{p(E,u)}{E}\,N_{0},$ (2)
is computed from the attenuated spectrum; $N_{0}$ is the incident photon fluence. Electronic read-out noise is added as correlated Gaussian noise. The result is a radiograph whose grain and contrast match clinical images.
The realism of DeepDRR has been validated in previous work [unberath2018deepdrr, gao2023synthetic], which demonstrated that models trained solely on DeepDRRs generalized effectively to real fluoroscopy images without any domain adaptation.
@@ III-B3 Generation of the X-Ray Dataset
The simulated X-ray dataset was generated from the same 150 CT scans described above, following the same patient-wise train/validation/test split. For each CT scan in the training and validation sets, 400 X-ray images were simulated to provide sufficient diversity in view direction. For the test set, a reduced sampling of 20 X-ray images per CT scan was used, resulting in a total of 600 test images. This lower sampling density was chosen to avoid excessive redundancy among similar projection angles and to comply with computational constraints of the Grand Challenge platform, which executes containerized inference on a per-case basis. The images were provided in binary Tag Image File Format (TIFF), with multi-label segmentation masks encoded as single-channel binary TIFF images to comply with the challenge platform requirements. Additionally, the encoding and decoding scripts for handling segmentation files were made available to participants to ensure consistency in data processing. The X-ray training set is publicly available in an open repository under a CC BY-NC-SA license [^3] PENGWIN X-ray dataset: https://doi.org/10.5281/zenodo.10913195..
III-C Assessment Method
III-C1 Evaluation Metrics
@@
For both the CT and X-ray tasks, segmentation performance was evaluated using intersection over union (IoU), the 95th percentile Hausdorff distance (HD95), and average symmetric surface distance (ASSD). IoU was calculated on solid masks and reflects the volumetric overlap between the predicted and ground-truth regions, serving as a general indicator of segmentation completeness. HD95 and ASSD were computed on surface contours to capture spatial discrepancies along fragment boundaries. While ASSD represents the average surface deviation and reflects typical alignment accuracy, HD95 quantifies the worst-case boundary error, excluding outliers. This is particularly important in clinical scenarios where large local inaccuracies near fracture lines or joint boundaries may compromise diagnosis or treatment planning. The adoption of these metrics aligns with recommendations from the Metrics Reloaded initiative and has been validated in prior benchmarking efforts [faghani2022mitigating].
Beyond algorithmic benchmarking, these metrics carry direct clinical relevance. Accurate estimation of fragment volume and displacement—reflected by IoU—is crucial for assessing fracture severity and planning fixation strategies. Surface-level metrics such as ASSD and HD95 are highly pertinent to the precision of virtual fracture reductions and implant alignment. Prior work has shown that segmentation errors propagate into increased inaccuracies in reduction pose estimation, underscoring the critical role of robust segmentation as a foundation for downstream surgical planning [liu2025preoperative].
IoU measures the overlap between the predicted segmentation mask $S_{p}$ and the ground truth mask $S_{g}$:
$\text{IoU}=\frac{|S_{p}\cap S_{g}|}{|S_{p}\cup S_{g}|}.$ (3)
HD95 quantifies the worst-case segmentation error by measuring the 95th percentile of the distances between the predicted and ground-truth segmentation contours:
@@
$\text{ASSD}=\frac{1}{|C_{p}|+|C_{g}|}(\sum_{x\in C_{p}}\inf_{y\in C_{g}}d(x,y)+\sum_{y\in C_{g}}\inf_{x\in C_{p}}d(x,y)),$ (5)
where distances are computed in both directions.
@@
In addition to accuracy metrics, we also report the container runtime as an indicator of efficiency. Note that this runtime includes not only the model inference time but also the overhead from container initialization, image loading, and result saving.
III-C2 Ranking Scheme
@@
    Metric calculation: For each individual test case, the accuracy metrics were computed for both fracture segmentation and anatomical segmentation.
@@
    Tie-breaking: In the event of a tie, the container execution time of the submission was used as a tie-breaker, with faster models receiving a higher rank.
IV Results
IV-A Participation
@@ IV-B Algorithm Summary
IV-B1 CT Segmentation
Figure 4: Different fracture segmentation pipelines in the CT task. (a) Boundary-core decomposition scheme adopted by MIC-DKFZ, MedApp-AGH, and Sano. (b) Primary-secondary labeling scheme adopted by SMILE and MedIG.
TABLE II: Summary of the top-performing participating algorithms for CT segmentation. Spatial augmentation includes random scaling, mirroring, and flipping. Intensity augmentation includes noising, blurring, and histogram adjustments.